# Cite Pretrain: Retrieval-Free Knowledge Attribution for Large Language Models

### v1

@@ 4.1 Passive Indexing
Identifier Format
: We use the natural document text title $t_{i}\in\mathcal{T}$ as the identifier, as titles encapsulate salient content and align with the model’s text-based learning paradigm. Our preliminary experiments show that using titles as document identifiers leads to better memorization compared to numerical IDs and other semantically structured alternatives (Appendix E.1). Additionally, titles are scalable as the textual space is vast and allows renaming to avoid duplication when necessary.
Identifier Placement
: Following prior work [15], we append $t_{i}$ at the end of $c_{i}$ during pretraining, forming inputs of the form $c_{i}\to t_{i}$. This mirrors downstream tasks where citations follow generated text, facilitating natural learning of citation patterns. As a baseline, we also tested inserting $t_{i}$ after each sentence within $c_{i}$ [15], but this reduced fluency and is used for comparison.
@@ 4.2 Active Indexing: Forward Augmentation
Entity Extraction
: For each document $d_{i}$, we extract a set of $N$ salient entities $E_{i}=\{e_{i1},\dots,e_{iN}\}$ using an auxiliary LLM, where $N$ controls the augmentation scale. Each $e_{ij}$ is a key concept or entity in $c_{i}$, serving as an anchor for question generation.
Question-Answer Pair Generation
: For each entity-document pair $(e_{ij},d_{i})$, an LLM generates a set of question-answer pairs $\{(q_{ijk},a_{ijk})\}_{k=1}^{K_{ij}}$, where $K_{ij}\geq 1$ is the number of pairs per entity-document pair. Each question $q_{ijk}\in\Sigma^{*}$ references $t_{i}$ and probes information related to $e_{ij}$ (e.g., who, what, where, why, how, if). Each answer $a_{ijk}\in\Sigma^{*}$ provides a detailed response based on $c_{i}$, containing facts from $S_{i}$. This creates a closed-book training signal that strengthens the mapping $t_{i}\to S_{i}$, encouraging the model to internalize and retrieve facts when prompted with $t_{i}$. We then post-process the noisy doc-IDs in the generated questions. See details in Appendix G.1.
@@ 4.3 Active Indexing: Backward Augmentation
Document Chunking and Indexing
: Each document $d_{i}$ is divided into a set of chunks $\mathcal{C}_{i}=\{c_{i1},\dots,c_{im_{i}}\}$, where each chunk $c_{ij}\in\Sigma^{*}$ contains $W$ words. The corpus-wide chunk set is $\mathcal{C}=\bigcup_{i=1}^{N}\mathcal{C}_{i}$. Chunks are indexed using retrieval methods (e.g., BM25), creating an index base $\mathcal{I}:\mathcal{C}\to\mathbb{R}^{k}$, where $\mathcal{I}(c_{ij})$ is the chunk’s representation.
Chunk Cluster Formation
: A chunk cluster $\mathcal{C}_{\ell}=\{c_{\ell 1},\dots,c_{\ell M_{\ell}}\}\subseteq\mathcal{C}$ is a set of related chunks from distinct documents. To form $\mathcal{C}_{\ell}$, we randomly sample $N$ seed chunks $\{c_{i1},\dots,c_{iN}\}$ from each document content $c_{i}$, where $N$ controls the augmentation scale. For each seed chunk $c_{ij}$, we retrieve $M$ relevant chunks $\{c_{\ell 1},\dots,c_{\ell M}\}$ from $\mathcal{I}$, where $M\sim\text{Uniform}(2,4)$ and each $c_{\ell m}$ belongs to a distinct document $d_{k}$, $k\neq i$.
Instruction-Answer Pair Generation
: For each chunk cluster $\mathcal{C}_{\ell}$, an LLM generates an instruction-answer pair $(q_{\ell},\mathcal{R}_{\ell})$, where $q_{\ell}\in\Sigma^{*}$ is an instruction requiring integration of information from $\mathcal{C}_{\ell}$, and =$\mathcal{R}_{\ell}=\{(s_{\ell k},C_{\ell k})\}_{k=1}^{m_{\ell}}$ is the response, with $s_{\ell k}$ a factual statement and $C_{\ell k}\subseteq\{t_{i}\mid c_{ij}\in\mathcal{C}_{\ell}\}$ the set of supporting titles. This aligns with downstream tasks where $g:q\to\{(s_{k},C_{k})\}$. To manage computational costs, we bootstrap a seed set of pairs using GPT-4.1-mini and fine-tune a Qwen-2.5-3B model to generate further augmentations. We post-filter the instance with invalid doc-IDs (around 5%). See details in Appendix G.2.
@@ 5.1 Experimental Details
We evaluate our approach using Qwen-2.5 models at both 7B and 3B scales, testing a range of methods to assess their effectiveness in enabling LLMs to cite from pretraining data. Evaluation is conducted across four QA benchmarks—ASQA, ELI5, RepliQA, and SciQAG (see Appendix A for dataset details). Results are reported in terms of both answer correctness and citation quality, measured by citation recall and citation precision. All methods follow a two-stage process: continual pretraining followed by instruction tuning. They differ only in their approach to continual pretraining. The Instruction-only method skips continual pretraining and proceeds directly to instruction tuning. We compare the following seven methods:
Instruction-only: Serves as a baseline. The base model is directly instruction-tuned without continual pretraining, enabling evaluation of citation behavior that relies solely on memorized pretraining knowledge.
@@
Repetition: Following prior work [15], document IDs are appended after each fact during continual pretraining to support fine-grained attribution, though frequent insertions may impair fluency.
@@
While data generation and continual pretraining add one-time overhead, our approach has no inference cost—unlike RAG methods, which incur ongoing retrieval and conditioning costs. See training details in Appendix C.
@@
Table 2: Main results on four question answering tasks. Acc = answer correctness, C-pr = citation precision, C-re = citation recall. Best results are bolded. We find that: (1) Active Indexing outperforms Passive Indexing; (2) Forward and Backward augmentation are complementary; (3) Larger models perform better.
@@ 5.2 Main Results
Passive Indexing Is Insufficient for Citation
Table 2 shows that Passive Indexing—simply appending document IDs and expecting implicit learning—is insufficient. While it offers modest gains over instruction tuning alone (e.g., 2.4 vs. 0.9 on RepliQA), citation precision remains low. Attaching doc-IDs closer to facts (Repeat) also fails to improve performance on realistic tasks, unlike prior work 15, likely due to their use of synthetic, rigidly structured data. In real-world settings, diverse and loosely aligned facts limit the effectiveness of sentence-level associations. Nonetheless, Repeat+ performs slightly better, likely due to its more granular ID attachment across different parts of the document, which enhances fluency and supports better generalization.
Active Indexing is Effective in Both Directions
Active Indexing consistently outperforms passive baselines across all tasks (Table 2). Backward indexing is more effective than forward, and combining both yields the largest gains (e.g., citation precision: 2.4 → 32.6), highlighting their complementarity. Scaling augmentation further could enhance performance (see Section 5.3). Notably, Active Indexing also boosts answer correctness, likely due to exposure to factual content in diverse formats [34, 35].
Model Size Matters
On the Qwen-2.5-3B model, we observe similar trends to the 7B variant; however, its absolute citation performance is substantially lower. For example, on SciQAG with active indexing, while the accuracy remains comparable, citation precision drops from 32.6 to 20.0, suggesting that model size plays a critical role in citation capability.
Pretrained Models Memorize Wikipedia Titles
On ASQA—a task based on the Wikipedia corpus—models achieve competitive citation performance even without indexing (e.g., 20.0 with instruction-tuning only vs. 24.1 with passive indexing). We hypothesize two main reasons: (1) large models memorize not only verbatim text during pretraining [23, 25] but also the corresponding Wikipedia titles; and (2) Wikipedia articles are both topically coherent and centered around well-defined entities, making it easier for models to associate content with document titles. Preliminary experiments on other Wikipedia-based tasks, such as TriviaQA and HotpotQA, show similar trends. See Appendix E.2 for details.
@@ 5.3 Why does Active Indexing work?
Diverse Fact Representations Help Citation
Figure 2: Scaling Curve of Combining Backward and Forward on RepliQA
Active Indexing generates diverse fact variants—through paraphrasing, composition, and interaction—all tied to the same document ID. This diversity helps the model generalize and reliably cite, improving both memorization and utilization. To evaluate how diversity impacts citation ability, we study how citation performance scales with the amount of augmented data. For efficiency, we focus on the RepliQA dataset, where we continually pretrain only on RepliQA corpus and then instruction-tune on its QA training set. We augment the RepliQA corpus: 66M tokens from Active Indexing Forward (AIF), 76M from Backward (AIB), and their combination (AI), totaling 16.5× the original size. From these, we subsample different proportions to construct a scaling curve. Figure 2 shows citation precision on RepliQA (y-axis) versus the fraction of augmented data used (x-axis). Key findings include: 1. Scaling helps: All variants—AIF, AIB, and AI—improve with more augmented data, even up to 16.5×. 2. Complementary directions: AIF alone is less effective, but combining AIF with AIB (i.e., AI) consistently outperforms either alone. Each point on the AI curve uses matched subsets of AIF and AIB, indicating additive and reinforcing effects. These results support that both the diversity and the quantity of augmentations play key roles in enabling accurate citation. 3. Room to grow: Citation performance continues to rise at our highest augmentation levels, suggesting the trend isn’t saturated and further gains are possible beyond our budgeted limits.
Actively Teaching Identifier Usage Is Crucial
While increasing the quantity and diversity of facts is important, it is not sufficient. The key advantage of active indexing lies in explicitly teaching the model to use document identifiers within QA-style contexts—directly aligning pretraining with the downstream citation objective. In contrast, passive indexing simply appends document identifiers to the corpus and replays it, hoping the model implicitly learns the association. We compare this against two passive indexing baselines: 1. More tokens (PI-Replay): Passive indexing with additional epochs. 2. More diversity (PI-SCP): Synthetic Continual Pretraining (SCP) [35], which paraphrases facts and introduces relational variants, but still lacks explicit QA-style training to use document IDs.
@@
Figure 3 shows citation precision on RepliQA (y-axis) as a function of the augmentation-to-original token ratio (x-axis). Passive indexing with replay fails to improve citation performance. It may even impair generalization, as the model overfits to shallow patterns. Passive Indexing with SCP fares better, supporting the role of diverse formulations, but remains far less token-efficient than active indexing. These findings highlight that active indexing’s strength goes beyond exposing the model to diverse facts. Its effectiveness comes from actively aligning model behavior with the task of attribution—by reinforcing identifier usage in QA, presenting compositional and contextual examples, and training the model to ground its answers in specific sources. This explicit supervision is critical to reliable citation performance.
@@
Past work has shown that stronger memorization does not gaurantee better generalization [36]. We observe a similar pattern in citation tasks. There is a crucial difference between how LLMs (1) memorize document identifiers during continual pretraining, and (2) generalize to use these identifiers when answering downstream questions. To probe this, we evaluate four setups that progressively shift from pure memorization to downstream usage: 1. FullDoc: Predicting the doc-ID given the full document. 2. PartialDoc: Predicting the doc-ID from a partial document segment. 3. GoldQA: Predicting the doc-ID given the question and gold answer. 4. ModelQA: Predicting the doc-ID given the question and the model’s generated answer. We reuse RepliQA 7B models and report Hit@1/10 accuracy. As shown in Table 3, all methods’ performance declines as tasks shift from memorization to generalization in downstream citations. Notably, more replay epochs (PassIdx-REP) improve memorization (FullDoc: 27.0 → 74.6) but hurt downstream generalization (ModelQA: 7.8 → 6.0), suggesting overfitting to shallow patterns. In contrast, Active Indexing bridges this gap by promoting both memorization and effective usage in QA, encouraging robust fact-to-source grounding.
@@
Table 3: Document ID Memorization and Generalization.
@@ 6 Conclusion
We demonstrate that large language models can attribute answers to their pretraining data without relying on test-time retrieval. To this end, we (1) introduce a benchmark for internal citation across both short- and long-form QA, and (2) propose Active Indexing, a continual pretraining strategy that teaches models to link content with document identifiers. Our experiments highlight three key findings: 1. Teaching beats hoping: Simply appending IDs (Passive Indexing) or replaying more tokens yields little benefit. In contrast, framing ID usage as a QA task (Active Indexing) boosts citation precision and recall by up to 32 points across four datasets, bridging memorization and utilization. 2. Complementary directions: Combining forward and backward augmentations yields stronger gains than either alone. 3. Scale helps: Citation performance improves with more augmented pretraining data. These results show that attribution can be embedded directly into model training, simplifying deployment and enabling more accountable language models. See Appendix H for a discussion of limitations and future directions.
Acknowledgments and Disclosure of Funding
Yukun Huang, Sanxing Chen, and Bhuwan Dhingra’s research is supported in part by the NSF award IIS-2211526. Jian Pei’s research is supported in part by the NSF DMS-2434666.
All opinions, findings, conclusions and recommendations in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.
Replication
The code and data will be released.

### v2

@@ 4.1 Passive Indexing
Identifier Format: We use the natural document text title $t_{i}\in\mathcal{T}$ as the identifier, as titles encapsulate salient content and align with the model’s text-based learning paradigm. Our preliminary experiments show that using titles as document identifiers leads to better memorization compared to numerical IDs and other semantically structured alternatives (Appendix E.1). Additionally, titles are scalable as the textual space is vast and allows renaming to avoid duplication when necessary.
Identifier Placement: Following prior work Khalifa et al. (2024), we append $t_{i}$ at the end of $c_{i}$ during pretraining, forming inputs of the form $c_{i}\to t_{i}$. This mirrors downstream tasks where citations follow generated text, facilitating natural learning of citation patterns. As a baseline, we also tested inserting $t_{i}$ after each sentence within $c_{i}$ (Khalifa et al., 2024), but this reduced fluency.
@@ 4.2 Active Indexing: Forward Augmentation
Entity Extraction: For each document $d_{i}$, we extract a set of $N$ salient entities $E_{i}=\{e_{i1},\dots,e_{iN}\}$ using an auxiliary LLM, where $N$ controls the augmentation scale. Each $e_{ij}$ is a key concept or entity in $c_{i}$, serving as an anchor for question generation.
Question-Answer Pair Generation: For each entity-document pair $(e_{ij},d_{i})$, an LLM generates a set of question-answer pairs $\{(q_{ijk},a_{ijk})\}_{k=1}^{K_{ij}}$, where $K_{ij}\geq 1$ is the number of pairs per entity-document pair. Each question $q_{ijk}\in\Sigma^{*}$ references $t_{i}$ and probes information related to $e_{ij}$ (e.g., who, what, where, why, how, if). Each answer $a_{ijk}\in\Sigma^{*}$ provides a detailed response based on $c_{i}$, containing facts from $S_{i}$. This creates a closed-book training signal that strengthens the mapping $t_{i}\to S_{i}$, encouraging the model to internalize and retrieve facts when prompted with $t_{i}$. We then post-process the noisy doc-IDs in the generated questions. See details in Appendix G.1.
@@ 4.3 Active Indexing: Backward Augmentation
Document Chunking and Indexing: Each document $d_{i}$ is divided into a set of chunks $\mathcal{C}_{i}=\{c_{i1},\dots,c_{im_{i}}\}$, where each chunk $c_{ij}\in\Sigma^{*}$ contains $W$ words. The corpus-wide chunk set is $\mathcal{C}=\bigcup_{i=1}^{N}\mathcal{C}_{i}$. Chunks are indexed using retrieval methods (e.g., BM25), creating an index base $\mathcal{I}:\mathcal{C}\to\mathbb{R}^{k}$, where $\mathcal{I}(c_{ij})$ is the chunk’s representation.
Chunk Cluster Formation: A chunk cluster $\mathcal{C}_{\ell}=\{c_{\ell 1},\dots,c_{\ell M_{\ell}}\}\subseteq\mathcal{C}$ is a set of related chunks from distinct documents. To form $\mathcal{C}_{\ell}$, we randomly sample $N$ seed chunks $\{c_{i1},\dots,c_{iN}\}$ from each document content $c_{i}$, where $N$ controls the augmentation scale. For each seed chunk $c_{ij}$, we retrieve $M$ relevant chunks $\{c_{\ell 1},\dots,c_{\ell M}\}$ from $\mathcal{I}$, where $M\sim\text{Uniform}(2,4)$ and each $c_{\ell m}$ belongs to a distinct document $d_{k}$, $k\neq i$.
Instruction-Answer Pair Generation: For each chunk cluster $\mathcal{C}_{\ell}$, an LLM generates an instruction-answer pair $(q_{\ell},\mathcal{R}_{\ell})$, where $q_{\ell}\in\Sigma^{*}$ is an instruction requiring integration of information from $\mathcal{C}_{\ell}$, and $\mathcal{R}_{\ell}=\{(s_{\ell k},C_{\ell k})\}_{k=1}^{m_{\ell}}$ is the response, with $s_{\ell k}$ a factual statement and $C_{\ell k}\subseteq\{t_{i}\mid c_{ij}\in\mathcal{C}_{\ell}\}$ the set of supporting titles. This aligns with downstream tasks where $g:q\to\{(s_{k},C_{k})\}$. To manage computational costs, we bootstrap a seed set of pairs using GPT-4.1-mini and fine-tune a Qwen-2.5-3B model to generate further augmentations. We post-filter the instance with invalid doc-IDs (around 5%). See details in Appendix G.2.
@@ 5.1 Experimental Details
We evaluate our approach using Qwen-2.5 3B and 7B models, testing different methods to assess their effectiveness in enabling LLMs to cite from pretraining data. Evaluation is conducted across four QA benchmarks—ASQA, ELI5, RepliQA, and SciQAG (see Appendix A for dataset details). Results are reported in terms of both answer correctness and citation quality, measured by citation recall and citation precision. All methods follow a two-stage process: continual pretraining followed by instruction tuning. They differ only in their approach to continual pretraining. The Instruction-only method skips continual pretraining and proceeds directly to instruction tuning. We compare:
Instruction-only: Serves as a baseline: the model is instruction-tuned without continual pretraining, testing citation based solely on pre-trained knowledge.
@@
Repetition: Following Khalifa et al. (2024), document IDs are appended after each fact to support fine-grained attribution, though frequent insertions may reduce fluency.
@@
[…]
Table 1: Main results on four QA datasets. Acc=answer correctness, C-Pr=citation precision, C-Re=citation recall. Best results are bolded. We find that: (1) Active Indexing outperforms Passive Indexing; (2) Forward and Backward are complementary; (3) Larger models help, but without Active Index, even proprietary LLMs still struggle with internal citation.
@@ 5.2 Main Results
Passive Indexing Is Insufficient for Citation. Table 1 shows that Passive Indexing—simply appending document IDs and expecting implicit learning—is insufficient. While it offers modest gains over instruction tuning alone (e.g., 2.4 vs. 0.9 on RepliQA), citation precision remains low. Attaching doc-IDs closer to facts (Repeat) also fails to improve performance on realistic tasks, unlike prior work Khalifa et al., 2024, likely due to their use of synthetic, rigidly structured data. In real-world settings, diverse and loosely aligned facts limit the effectiveness of sentence-level associations. Nonetheless, Repeat+ performs slightly better, likely due to its more granular ID attachment across different parts of the document, which enhances fluency and supports better generalization.
Active Indexing is Effective in Both Directions. Active Indexing consistently outperforms passive baselines across all tasks (Table 1). Backward indexing is more effective than forward, and combining both yields the largest gains (e.g., citation precision: 2.4 → 32.6), highlighting their complementarity. Scaling augmentation further could enhance performance (see §5.3). Notably, Active Indexing also boosts answer correctness, likely due to exposure to factual content in diverse formats (Allen-Zhu & Li, 2025; Yang et al., 2025).
Model Size Matters. On the Qwen-2.5-3B model, we observe similar trends to its 7B counterpart, but with substantially lower absolute citation performance. For example, on SciQAG with Active Indexing, accuracy remains comparable, yet citation precision drops from 32.6 to 20.0—highlighting the importance of model size in citation capability. However, strong models like GPT-4.1 struggle with internal citation without targeted training: while they can occasionally predict Wikipedia titles, they often fail to predict titles for arXiv papers. This limitation is especially pronounced on tasks like ELI5, which depend on Common Crawl sources that lack document identifiers during pretraining, rendering internal citation infeasible.
Pretrained Models Memorize Wikipedia Titles. On ASQA (Wikipedia-based task), models achieve competitive citation performance even without indexing (e.g., 20.0 with instruction-tuning only vs. 24.1 with passive indexing). We attribute this to two main factors: (1) large models memorize not just verbatim text but also Wikipedia titles during pretraining (Weller et al., 2024; Zhang et al., 2025b), and (2) Wikipedia titles are typically entities, making it easier for models to “shortcut” by predicting the entity as the title. Preliminary results on other Wikipedia-based tasks, such as TriviaQA and HotpotQA, show similar patterns. See Appendix E.2 for details.
@@ 5.3 Why does Active Indexing work?
Facts Variation and Active Supervision Are Both Crucial for Reliable Citations
@@
To identify what drives reliable internal citation, we study how fact variation (i.e., presenting the same facts in different ways) and active supervision affect model performance. Using the RepliQA dataset, we scale the amount of augmented data and evaluate how citation performance varies with the augmentation ratio (Figure 2). We compare three approaches: (1) Passive Indexing with Replay, which simply repeats the same facts and document identifiers with more epochs; (2) Passive Indexing with Synthetic Continual Pretraining (Yang et al., 2025) (PI-SCP), which paraphrases facts and introduces relational variants, but still lacks explicit QA-style training to use document IDs. and (3) Active Indexing, which not only diversifies fact formulations but also explicitly trains the model to use document identifiers in QA-style contexts. Our findings are threefold: 1. Fact Variation Helps: Citation performance improves as Active Indexing increases both the scale and diversity of fact presentations. In contrast, simple token replay provides no benefit—and can even degrade performance due to overfitting. 2. Active Supervision is Crucial: Although PI-SCP introduces diverse phrasings of the same facts, it still lags behind active indexing. This indicates that diversity alone is insufficient—explicitly teaching the model to use document identifiers in context is essential for robust citation. 3. Room to grow: Citation performance continues to improve at the highest augmentation levels tested ($16\times$), suggesting that further gains are possible beyond our current limits.
@@
Past work has shown that stronger memorization does not gaurantee better generalization (Wang et al., 2025b). We observe a similar pattern in citation tasks. There is a crucial difference between how LLMs (1) memorize document identifiers during continual pretraining, and (2) generalize to use these identifiers when answering downstream questions. To probe this, we evaluate four setups that progressively shift from pure memorization to downstream usage: 1. FullDoc: Predicting the doc-ID given the full document. 2. PartialDoc: Predicting the doc-ID from a partial document segment. 3. GoldQA: Predicting the doc-ID given the question and gold answer. 4. ModelQA: Predicting the doc-ID given the question and the model’s generated answer. We reuse RepliQA 7B models and report Hit@1/10 accuracy. As shown in Table 2, all methods’ performance declines as tasks shift from memorization to generalization in downstream citations. Notably, more replay epochs (PassIdx-REP) improve memorization (FullDoc: 27.0 → 74.6) but hurt downstream generalization (ModelQA: 7.8 → 6.0), suggesting overfitting to shallow patterns. In contrast, Active Indexing bridges this gap by promoting both memorization and effective usage in QA, encouraging robust fact-to-source grounding.
Table 2: Document ID Memorization and Generalization.
@@
5.4 Internal vs. External Citations
Figure 3: Performance of internal, external, and hybrid citations across retrieval quality (0=sparse retrieval, 1=dense retrieval). Internal only excels under poor retrieval, external only under strong retrieval, while hybrids generally perform best regardless of retrieval quality, with room to improve.
We compare internal and external citation methods to understand their respective limitations and the potential for synergy. External citations allow models to access information beyond their memorized knowledge, while internal citations serve as a robust fallback when retrieval fails. This trade-off becomes especially important when retrieval quality varies. To explore this, we compare:
Internal Only: Generate internal citations without retrieval using Active Index.
External Only: Generate citations via RAG with 3-shot examples and top-5 retrieved documents, following the best-performing setup in Gao et al. (2023c).
Hybrid Joint: Instruction-tuning ActiveIdx to consume retrieved documents, generating both internal and external citations after an initial assessment of document sufficiency during inference.
Hybrid Fallback: A pipeline that first attempts RAG, falling back to Hybrid Joint if retrieved documents are deemed insufficient by the model.
Hybrid Oracle: A conceptual upper-bound that selects the better output from the Internal and External Only methods for each example. See more experiments details in Appendix C.2.
We test these strategies across a spectrum of retrieval quality, which we simulate by interpolating between sparse retrieval (BM25; lower-quality in our setup) and dense retrieval (Lin et al., 2021; high-quality). As shown in Figure 3, when retrieval is poor, methods incorporating internal citations (Internal Only and Hybrid variants) significantly outperform External Only, highlighting the importance of internal fallback under noisy retrieval. As retrieval improves, external outperforms internal, showing that reliable external evidence can complement internal knowledge. Hybrid Approach, taking advantages of both, perform generally best regardless of retrieval quality. Notably, a performance gap remains between our best Hybrid method and the Hybrid Oracle, pointing to headroom for more effective integration strategies that better reconcile retrieved and memorized knowledge.
@@ 6 Conclusion
We show that large language models can attribute answers to their pretraining data without relying on test-time retrieval. We (1) introduce CitePretrainBench for internal citation across short- and long-form QA, and (2) propose Active Indexing, a continual pretraining strategy that teaches models to link content with document identifiers. Key findings include: 1. Teaching beats hoping: Active Indexing, which frames citation as real tasks, improves precision and recall by up to 32 points—outperforming passive approaches. 2. Complementary directions: Forward and backward augmentations are most effective when combined. 3. Scale helps: Performance improves consistently with more augmented data. 4. Internal and external are complementary: Combining both enhances robustness to retrieval quality. See Appendix H for limitations and future directions.
Reproducibility
We are committed to ensuring the reproducibility of our results. We describe details of our training procedures (§4,) and implementations (§5.1, Appendix C). For our evaluation, we describe details of datasets (§3.2), metrics (§3.3, §3.4) and processing steps (Appendix A). Code and data will be released upon publication.